# Tight Robustness Certificates and Wasserstein Distributional Attacks for Deep Neural Networks

### v1

@@ A.3 Technical Proofs
Cross-Entropy Loss: The difference in loss is $\Delta\ell_{CE}=\ell_{CE}(\theta(\tilde{x}),e_{k^{\star}})-\ell_{CE}(\theta(x^{\star}),e_{k^{\star}})$. Using the property $\ell_{CE}(z,e_{k^{\star}})=-(z_{k^{\star}}-\log\sum_{k}e^{z_{k}})$, the loss difference is:
@@
DLR Loss: The difference in DLR loss is $\Delta\ell_{DLR}=\ell_{DLR}(\theta(\tilde{x}),k^{\star})-\ell_{DLR}(\theta(x^{\star}),k^{\star})$.
@@
Appendix B Additional result
Figure 7: Comparison of step size effect on model Cui et al. (2024) under $r_{\infty}$ perturbation ($\epsilon=8/255$).
Figure 7 illustrates the convergence behavior of Cui et al. (2024) in terms of robust accuracy. Larger step sizes (0.1, 0.5) lead to higher final accuracy, whereas a smaller step size of 0.02 results in the lowest robust accuracy, indicating the most effective attack.
Figure 8: Comparison of robust accuracy with varying $\kappa$
Figure 8 presents the robust accuracy across different values of $\kappa$. Among the tested settings, $\kappa=2$ consistently produces the lowest robust accuracy for all models (Cui et al. (2024), Wang et al. (2023), and Sehwag et al. (2022)).
Appendix C Usage of Large Language Models (LLMs)
We used ChatGPT solely for revising the writing of the paper. Note that revision here strictly means enhancing the clarity and readability of the text (e.g., fixing typos or constructing latex tables), and not for any other purposes.

### v2

@@ Appendix C Auxiliary Results
Cross-Entropy Loss.
The difference in loss is $\Delta\ell_{CE}=\ell_{CE}(\theta(\tilde{x}),e_{k^{\star}})-\ell_{CE}(\theta(x^{\star}),e_{k^{\star}})$. Using the property $\ell_{CE}(z,e_{k^{\star}})=-(z_{k^{\star}}-\log\sum_{k}e^{z_{k}})$, the loss difference is:
@@
DLR Loss.
The difference in DLR loss is $\Delta\ell_{DLR}=\ell_{DLR}(\theta(\tilde{x}),k^{\star})-\ell_{DLR}(\theta(x^{\star}),k^{\star})$.
@@
Corollary C.2 (Feasibility).
Let $\mathbb{P}_{N}=\sum_{i=1}^{N}\mu_{i}\bm{\delta}_{(x^{(i)},y^{(i)})}$ and $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$. For $\alpha_{i}\in[0,1]$ and points $\tilde{x}^{(i)}\in\mathcal{X}$, define
$\mathbb{P}_{\rm adv}\triangleq\sum_{i=1}^{N}\mu_{i}\Bigl((1-\alpha_{i})\bm{\delta}_{(x^{(i)},y^{(i)})}+\alpha_{i}\bm{\delta}_{(\tilde{x}^{(i)},y^{(i)})}\Bigr).$
Then for any $p\in[1,\infty)$,
$\mathcal{W}_{d,p}(\mathbb{P}_{\rm adv},\mathbb{P}_{N})^{p}\leq\sum_{i=1}^{N}\mu_{i}\alpha_{i}\|\tilde{x}^{(i)}-x^{(i)}\|_{r}^{p}.$
In particular, if $\sum_{i=1}^{N}\mu_{i}\alpha_{i}\|\tilde{x}^{(i)}-x^{(i)}\|_{r}^{p}\leq\epsilon^{p}$, then $\mathbb{P}_{\rm adv}\in\Omega_{p}(\epsilon)=\{\mathbb{P}:\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\}$.
Proof.
Use the explicit coupling that leaves $(1-\alpha_{i})$ mass of each atom fixed and transports $\alpha_{i}$ mass from $(\tilde{x}^{(i)},y^{(i)})$ to $(x^{(i)},y^{(i)})$. ∎
Appendix D WDA Study
The $\kappa$ parameter.
Figure 6 illustrates how the robust accuracy of WDA varies as the parameter $\kappa$ increases. Across all models, raising $\kappa$ beyond 1 generally leads to a noticeable drop in robust accuracy. Specifically, for the models from Figure 6(a) in $r_{2}$, $\epsilon=0.5$ settings, the best performance is observed at $\kappa=3$, whereas for Ding et al. (2020); Cui et al. (2024); Wang et al. (2023); Sehwag et al. (2022) in Figure 6(b), the lowest robust accuracy occurs at $\kappa=2$. These results indicate that increasing $\kappa$ can weaken model robustness, with the precise $\kappa$ that produces the largest drop depending on the architecture.
(a) CIFAR-10 – $r_{2}$, $\epsilon=0.5$.
(b) CIFAR-10 – $r_{\infty}$, $\epsilon=8/255$.
Figure 6: Analysis of robust accuracy across different $\kappa$ scales.
The step size parameter.
In Figure 7(a) ($r_{2},\epsilon=0.5$), we observe that step sizes significantly larger than the budget (e.g., 1.0 and 0.5) cause the fastest initial drop in accuracy but risk overshooting the optimal perturbation. In contrast, a moderate step size of 0.2 provides the most stable convergence to the lowest robust accuracy. Very small step sizes (0.01, 0.005) fail to explore the $\epsilon$-ball effectively within the given iterations, resulting in an underestimation of the attack’s potency. Similarly, in Figure 7(b) ($r_{\infty},\epsilon=8/255\approx 0.031$), the choice of step size is critical for convergence. A step size of 0.02 (approximately $2/3\epsilon$) yields the most potent attack. Step sizes significantly larger than $\epsilon$ (0.1, 0.5) lead to higher final robust accuracy due to oscillation around the boundary of the $L_{\infty}$ ball, while the smallest step size (0.005) suffers from insufficient progress, failing to reach the floor within 20 iterations.
(a) CIFAR-10 – $r_{2}$, $\epsilon=0.5$.
(b) CIFAR-10 – $r_{\infty}$, $\epsilon=8/255$.
Figure 7: Analysis of convergence and step size effects across different perturbation types. (a) Method from Wang et al. (2023); (b) Method from Cui et al. (2024).
Appendix E WDA++ Study
While the standard WDA fixes a global scalar $\kappa$ for all samples, WDA++ adopts an adaptive strategy to allocate the transport budget more efficiently. The core intuition is to assign higher transport mass (larger $1/\kappa_{i}$) to samples that are geometrically closer to the decision boundary, as these require less “cost” to flip. We report the classification accuracy on the constructed distribution $\mathbb{P}_{\mathrm{adv}}$ as
$\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\times\bm{1}_{\{h_{\theta}(x_{i})=y_{i}\}}+\frac{1}{\kappa_{i}}\times\bm{1}_{\{h_{\theta}(x^{\mathrm{adv}}_{i})=y_{i}\}},$
where $h_{\theta}(x)=\arg\max_{k}\theta_{k}(x)$ is the class predictor. The practical implementation involves three specific algorithmic components to ensure efficiency and optimality under the budget constraint. These are detailed below and correspond to the steps in Algorithm 2.
Algorithm 2 WDA++
Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\frac{1}{N}\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, Wasserstein order $p\in\{1,2\}$, step size $\alpha>0$, srchiter, maxiter, and topK
Outputs: adversary distribution $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,p}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$, where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
Initialize: dual-norm maximizer $\mathcal{M}$ equation 1
for $i=1$ to NN do
  if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{0})\neq k$ then $X_{\mathrm{adv}}^{(i)}\leftarrow X^{(i)}$, $d_{i}\leftarrow 0$ continue
  $x_{0}\leftarrow X^{(i)}$
  $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k\in\{1,\dots,K\}$
  $\mathcal{J}_{i}\leftarrow\arg\max_{\begin{subarray}{c}\mathcal{J}\subseteq\{1,\dots,K\}\setminus\{k\}\\ |\mathcal{J}|=\texttt{topK}\end{subarray}}\sum_{j\in\mathcal{J}}\theta_{j}(x_{0})$
  for iter =0=0 to maxiter do
   $g_{j}\leftarrow\nabla_{x}\theta(x_{\mathrm{iter}})^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}_{i}$
   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}_{i}$
   $\varphi_{j}\leftarrow x_{\mathrm{iter}}+\alpha u_{j}$ for $j\in\mathcal{J}_{i}$
   $j^{*}\leftarrow\arg\max_{j\in\mathcal{J}_{i}}\left(\theta_{j}(\varphi_{j})-\theta_{k}(\varphi_{j})\right)$
   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
   if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{\mathrm{iter+1}})\neq k$ then $x_{\mathrm{pre}}^{(i)}\leftarrow x_{\mathrm{iter}}$, $x_{\mathrm{flip}}^{(i)}\leftarrow x_{\mathrm{iter+1}}$ break
  end for
  if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{\mathrm{iter+1}})=k$ then $X_{\mathrm{adv}}^{(i)}\leftarrow X^{(i)}$, $d_{i}\leftarrow+\infty$ continue
  $\Delta\leftarrow x_{\mathrm{flip}}^{(i)}-x_{\mathrm{pre}}^{(i)}$, $t_{\ell}\leftarrow 0$, $t_{u}\leftarrow 1$
  for $j=1$ to srchiter do
   $t\leftarrow(t_{\ell}+t_{u})/2$
   $\tilde{x}\leftarrow x_{\mathrm{pre}}^{(i)}+t\Delta$
   if $\arg\max_{c}\theta_{c}(\tilde{x})\neq k$ then $t_{u}\leftarrow t$ else $t_{\ell}\leftarrow t$
  end for
  $X_{\mathrm{adv}}^{(i)}\leftarrow x_{\mathrm{pre}}^{(i)}+t_{u}\Delta$
  $d_{i}\leftarrow\|X_{\mathrm{adv}}^{(i)}-X^{(i)}\|_{r}$
end for
$\mathcal{I}\leftarrow\arg\mathrm{sort}\left(\left\{d_{i}:i\in\{1,\dots,N\}\right\}\right)$
$\kappa_{i}\leftarrow+\infty$ for $i\in\{1,\dots,N\}$
$B\leftarrow\epsilon^{p}$
for $i\in\mathcal{I}$ do
  if $B\leq 0$ then break
  $\kappa_{i}\leftarrow{1}/{\min\left\{1,\ {NB}/{d_{i}^{p}}\right\}}$
  $B\leftarrow B-\frac{1}{N}\frac{1}{\kappa_{i}}d_{i}^{p}$
end for
$\mathbb{P}_{\mathrm{adv}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa_{i}}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}$
return $\mathbb{P}_{\mathrm{adv}}$
Direction search via Top-K heuristic.
Evaluating the gradient for all K1K-1 target classes can be computationally expensive for datasets with many classes (e.g., CIFAR-100, ImageNet). To accelerate the attack, we employ a topK heuristic. We restrict the search for the optimal adversarial direction to the top-$K$ classes with the highest logits (excluding the correct class). In our experiments, we set $\texttt{topK}=20$ for ImageNet, $\texttt{topK}=10$ for CIFAR-100 and $\texttt{topK}=5$ for CIFAR-10.
Boundary estimation via binary search.
To compute the minimal flip cost, we cannot rely solely on the gradient step, as it may overshoot or undershoot the decision boundary. Instead, we perform a binary search along the direction of the successful perturbation vector $\Delta=x_{\mathrm{flip}}^{(i)}-x_{\mathrm{pre}}^{(i)}$. This refinement step, controlled by $\texttt{srchiter}=10$, gives $d_{i}$.
Budget allocation.
WDA++ adaptively allocates the global Wasserstein budget $\epsilon$ by treating the attack as a Knapsack-like optimization problem: maximizing misclassified samples subject to $\sum\mu_{i}d_{i}^{p}/\kappa_{i}\leq\epsilon^{p}$. We adopt an optimal greedy strategy that calculates the flip cost $d_{i}$ for each sample, sorts them in ascending order, and sequentially assigns maximum weight (up to $\kappa_{i}=1$) to the “cheapest” samples until the budget $B=\epsilon^{p}$ is exhausted. This prioritizes the weakest links in the model to maximize the error rate for a given distributional radius.